James, Keller, and Rivers have the following capital balances; P48,000, P70,000 and P90,000 respectively. Because of a cash shortage James invests an additional P12,000 on June 1st. Each partner withdraws P1,000 per month. James, Keller, and Rivers receive a salary of P13,000, P15,000 and P20,000, respectively, for work done during the year. Each partner receives interest of 8% on their weighted average capital balance without regard to normal drawings. Any remaining profits are split 20%, 30%, and 50% respectively. The net income for the year is P30,000.
1
Expert's answer
2020-06-08T12:04:58-0400
  1. Interest of capital

James

"0.08\times48000+0.08\times12000\times\frac{7}{12}=4400"

Keller

"0.08\times7000=5600"

Rivers

"0.08\times90000=7200"

2.Loss

4400+5600+7200+13000+15000+20000=65 200

65 200-30000-35 200

"35 200\times0.2=7040"

"35200\times0.3=10560"

"35 200\times0.5=17 600"

2.capital balances on end:

"James=48000+12000+13000+4400-1000\times12-7040=58 360"

"Keller=70 000+15000+5600-1000\times12-10560=68 040"

"Rivers=90000+20000+7200-1000\times12-17060=87 600"